1a. The drying time of a type of paint under specified test condition is known to...

1a.

The drying time of a type of paint under specified test condition is known to be normally distributed with an assumed mean value of 75 and a standard deviation of 9. Test the hypothesis H0: µ=75 vs. Ha: µ<75 from a sample of 25 observations. For a test procedure with α= 0.002 what is the probability of making a type 2 error when the true mean is 70, ß(70) (to 4 decimals)?

1b.

The drying time of a type of paint under specified test condition is known to be normally distributed with an assumed mean value of 75 and a standard deviation of 9. Test the hypothesis H0: µ=75 vs. Ha: µ<75 from a sample of 25 observations.  If the test procedure for α = 0.002 is used, what n is necessary to ensure that ß(70) = 0.01?
Answer #1

1a_

for 0.002 level with left tail test , critical z= -2.878 (from excel:normsinv(0.002)
sample size    n= 25
std deviation σ= 9.00
std error ='σx=σ/√n=9/√25= 1.800
rejection region: Xbar <=μ+Zα*σx or Xbar<= 75-2.87816174*1.8= 69.8193
P(Type II error) =P(Xbar>69.819|μ=70)=P(Z>(69.8193-70)/1.8)=P(Z>-0.1)=0.5398

1b)

Hypothesized mean μo= 75
true mean μa= 70
std deviation σ= 9.00
0.002 level critical Z= 2.878
0.01 level critical Zβ= 2.326
n=(Zα/2+Zβ)2σ2/(μoa)2= 88